OpenStudy (anonymous):

find the locus of points P(x, y) which moves so that the area of the triangle with vertices formed by the points P1(3, 5), P2(5, 8) and P has an area of 10 square units.

OpenStudy (anonymous):

 then A = 0.5*b*h

OpenStudy (anonymous):

you have to use distance formulae

OpenStudy (anonymous):

yeah figure that much, i'm just confused with the area thing >,<

OpenStudy (anonymous):

use area equation to get an equation interms of x and y

OpenStudy (anonymous):

should i assume that the base should be the two given points?

OpenStudy (anonymous):

\[10=0.5\times b\times h \] yes. from the diagram, for that fixed base, the height is fixed too.

OpenStudy (anonymous):

you'd get a number for b \[h^2={25\over b^2}\] equate the two H62 equations.

OpenStudy (anonymous):

ok i'll try it thanks

OpenStudy (anonymous):

um where did you get the 25?

OpenStudy (helder_edwin):

do you know a little about determinants?

OpenStudy (anonymous):

you can also use Heron's formula as follows: \[ 4A^2=|P_1P_2|^2|P_1P|^2-(P_1P_2-P_1P)^2 \]

OpenStudy (anonymous):

sorry i don't know anything about determinants or i don't remember it, sorry

OpenStudy (anonymous):

sorry. (10/0.5)^2=400 mistake..

OpenStudy (anonymous):

not 25

OpenStudy (anonymous):

oh its ok, your helping afterall, i don't have the right to complain^_^

OpenStudy (anonymous):

yes determinants.. \[ A = {1\over2}\left|\begin{matrix}x_0&y_0&1\\x_1&y_1&1\\x_2&y_2&1\end{matrix}\right| \]

OpenStudy (anonymous):

way easier this way

OpenStudy (anonymous):

so, \[ 2\times10=\left|\begin{matrix} x&y&1\\ 3&5&1\\ 5&8&1 \end{matrix}\right| \]

OpenStudy (anonymous):

oh that, so i'll just plug the points in there then copy the first two columns or not?

OpenStudy (anonymous):

you have to expand the determinant now

OpenStudy (anonymous):

ok try to figure out how to expand determinants, thanks